OpenStudy (anonymous):

Solve the equation. Check the solution. -4/x+1 = -1/x+5

OpenStudy (iambatman):

\[\frac{ -4 }{ x+1 }=-\frac{ 1 }{ x+5 }\] is it this?

OpenStudy (anonymous):

Yes

OpenStudy (iambatman):

You could "cross multiply" if you like and then solve for x

OpenStudy (iambatman):

\[\frac{ -4 }{ x+1 }=-\frac{ 1 }{ x+5 } \implies -4(x+5)=-1(x+1)\]

OpenStudy (anonymous):

No idea how to do any of that lol

OpenStudy (anonymous):

okay

OpenStudy (iambatman):

Now distribute and solve for x, can you do that

OpenStudy (anonymous):

nope I suck at math

OpenStudy (iambatman):

 multiply through what do you get?

OpenStudy (anonymous):

-20?

OpenStudy (iambatman):

No, you have to first multiply by x and then 5 so we get \[-4(x+5) = -4x-20\] does that make sense?

OpenStudy (anonymous):

okay yes i guess.

OpenStudy (iambatman):

Try \[-1(x+1)\] what do you get

OpenStudy (anonymous):

-x - 1

OpenStudy (anonymous):

right?

OpenStudy (iambatman):

Perfect!

OpenStudy (iambatman):

We have \[-4x-20=-x-1\] now we can collect "like terms" so the one's with the variables go together and then the ones which are just number are on its own.

OpenStudy (anonymous):

okay

OpenStudy (iambatman):

How do we do that?

OpenStudy (anonymous):

I don't know...

OpenStudy (iambatman):

Lets move all the terms with x on the left and the numbers on right, so -4x is already on the left, but notice the -x isn't there, so lets maybe add x on both sides and see what happens, \[-4x-20 \color{red}{+x}=-x-1 \color{red}{+x}\]

OpenStudy (iambatman):

Notice it gets cancelled on the right and we're left with the x on the right and we haven't really changed anything other than moving it to the other side :)

OpenStudy (anonymous):

okay! so what would i do now to figure out my answer?

OpenStudy (iambatman):

\[-4x+x-20=-1\] right now we're solving for x, what should we do next?